OpenStudy (anonymous):

Does anybody know how to solve inequalities with shown work? I undertand how,but these 2 problems confused me. Can you please help? 17-(4k-2) is greater than or equal to 2(k+3) 2n-3(n+3) is less than or equal to 14 Everytime i try these two problems i get a different answer. Can you please show step by step how to solve these. Thanks.

OpenStudy (anonymous):

Show us what you got, we'll point out any mistake.

OpenStudy (anonymous):

ok. for the first one i got k is less than or equal to 1.5 17-1(4k-2)is greater than or equal to2k+6 17-4k-2 is greater than or equal to 2k+6 17-17-4k-2+2 2k+6-17+2 -4k-2k -9 -6k divided by negative six -9 divided by -6 k=1.5

OpenStudy (phi):

notice that -1(4k-2) is -4k + 2 (not -2) 2nd, I would add +4k to both sides , to move the k's to the right side and move the pure numbers to the left side

OpenStudy (phi):

and remember this is >= not just =

OpenStudy (anonymous):

\[2n-3(n+3)\le14\] 2n-3n+6 14 -1n+6 14 -1n=6-6 14-6 _1n 8 divide each side by negative one and reverse the inequality \[n \ge -8\]

OpenStudy (phi):

you are making mistakes when you distribute. do it in smaller steps -3(n+3) = -3*n + -3*3 = -3n-9

OpenStudy (anonymous):

thank You

OpenStudy (anonymous):

So for the first one should i get \[k \le 1.5\]

OpenStudy (anonymous):

and for the second one \[n \ge -23\]

OpenStudy (phi):

2nd one is good. 1st one: 17-(4k-2) ≥ 2(k+3) 17-4k+2 ≥ 2k+6 19 -6 ≥ 2k+4k 13 ≥ 6k k≤ 13/6

OpenStudy (anonymous):

Thanks. So for the first one it wont be 13 divided by 6 it will be 13 over 6. Right?

OpenStudy (phi):

it is the same thing. but 13/6 is easier to type than 2.1666.... (6 repeating)

OpenStudy (phi):

you could also write \( 2\frac{1}{6} \)